# Mitigating Polarization in Recommender Systems via Network-aware Feedback Optimization

### v1

@@ 4.1 Opinion Dynamics and Clicking Behaviour
As a proof of concept, we consider a platform with $n=15$ users, and randomly generate the adjacency matrix as $A\sim\mathcal{U}[0,1]^{n\times n}$, $A\mathbf{1}_{n}\leq\mathbf{1}_{n}$ and their initial opinions as $x^{0}\sim\mathcal{U}[-1,1]^{n}$. Moreover, we set $\Gamma_{d}=0.5\mathrm{diag}[|x^{0}|]$. This choice of $\Gamma_{d}$ indicates that the users who hold initial extreme opinions tend to be more stubborn. The external influence $d$ is modelled as a prejudice term defined as the initial opinion $x^{0}$. Further, $\Gamma_{p}$ is drawn from $\mathcal{D}_{[10^{-2},0.5]}^{n}$ so that all users are affected by the recommendations. To incorporate diversity in clicking behaviours, we randomly assign $8$ users to follow clicking behaviour $C_{A}$ and the remaining ones follow $C_{B}$.
@@ 4.2 Experimental Setting and Hyperparameter Selection
In the online phase, we choose $\sigma_{\text{pe}}=0.07$ in (16) for the first $N/5$ time steps to facilitate sensitivity estimation and network learning, and $\sigma_{\text{pe}}=0$ afterwards to favor accurate recommendations. In the Kalman filter (3.2.1), we initialize ^$\hat{\ell}^{0}=0.5{\rm vec}(I_{n})$ and $\Sigma^{0}=I_{n^{2}}$. For gradient estimation in (14)–(15), we set the smoothing parameter $\mu=0.1$. While tuning the process noise standard deviation, we set the tuning parameter $M$ in (21) as $M=10$. Further, we start with neutral recommendations, i.e., $p^{0}=0_{n}$, and consider a periodic trigger with period $T=60$. For the polarization cost $\varphi^{\rm pol}$ in (4), we set $\epsilon_{1}=-0.5$ and $\epsilon_{2}=0.5$. All the simulations are conducted for $N=2.5\cdot 10^{4}$ time instances and over $50$ Monte-Carlo trials, with each simulation taking around 2 seconds. Since clicks are stochastic, the sequence of clicks varies across each trial. Moreover, the dither signal $w_{\rm pe}$ sequence varies across each trial, thus leading to different sensitivity estimates.
@@ 4.3 Architecture Comparisons
In Table 2, we summarize all the methods used for comparison and their attributes. In particular, for the oracle (method $M_{1}$), we employ the standard projected-gradient controller (6), wherein opinions can be directly measured, and the sensitivity and users’ clicking behavior is known.
@@
Figure 3 shows that convergence is empirically established in all the methods through the fixed-point residual mapping norm $\lVert\mathcal{G}(p^{k})\rVert^{2}$. In particular, for methods that involve sensitivity estimation (i.e., $M_{2}-M_{4}$), we observe a higher fluctuation for the first $N/5$ time instances due to the effect of the dither signal $\sigma_{\text{pe}}$. Our proposed method offers slightly inferior performances compared to the other OFO methods. However, it must be taken into account that method $M_{4}$ is agnostic about the analytical form of $g(p,x)$, and is completely data-driven.
@@
Figure 4: Illustration of sensitivity estimation error in methods $M_{2}-M_{4}$. The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
@@
4.4 Impact of Network Aware Recommendations
@@
We showcase the impact of providing network-aware recommendations compared to individual, decoupled recommendations. To achieve this, we compare our approach with two network-agnostic methods: (i) the recommender system described in [10], and (ii) a naive version of our OFO-based recommender design.
@@
The design in [10] provides recommendations with exclusively extreme positions (i.e., $p_{i}=\pm 1$) to each user independently. The position is calculated based on the user’s individual clicking history. Specifically, if a user more frequently clicks on $+1$ (or $-1$) positions, the system will provide $+1$ (or $-1$) positions with high probability. Conversely, if the user shows an equal likelihood of clicking on both $+1$ and $-1$ positions, the recommendations will be random.
@@
The latter is a network-agnostic version of our recommender system design summarized in Algorithm 1. Specifically, for each user $i\in[n]$, we use a local ANN for opinion estimation that is trained using local positions and acceptance ratio only, i.e., $\hat{x}_{i}=\hat{\beta}_{i}(y_{i},p_{i})$ rather than $\hat{\beta}_{i}(y_{i},p)$. Further, we do not carry out sensitivity estimation and instead use a constant diagonal sensitivity $\hat{H}\in\mathcal{D}_{(0,0.5]}^{n}$ in (17). The rationale behind this design choice is that if the network did not contribute to engagement maximization and/or polarization mitigation, a diagonal $\hat{H}$ would yield the same performance as our network-aware approach.
@@
We compare the performance of our recommender system with the other decoupled methods in terms of final mean CTR $J:=\sum_{k=N-T+1}^{N}\lVert c^{k}\rVert_{1}/(nT)$ and final polarization cost $\varphi^{\rm pol}(x^{N})$ in Figure 7 and the fixed point residual mapping in Figure 8. In Figure 7, we observe that our proposed algorithm achieves a slightly higher user engagement compared to its naive implementation and results in a lower polarization cost. Thus, network-aware recommendations effectively reduce polarization while maintaining high user engagement. This performance improvement is further illustrated in Figure 8, which depicts the temporal evolution of the fixed-point residuals. This result is unsurprising since polarization is a global network metric, whereas engagement maximization can be addressed individually by tailoring recommended content for each user.
@@
The recommender design in [10], which employs only extreme recommendations, naturally leads to higher polarization costs, as shown in Figure 7(b). Furthermore, it results in lower user engagement compared to both our algorithm and the naive method. This is because some users in the network are assigned clicking behavior $C_{B}$, for which extreme recommendations do not necessarily increase engagement. Also, the setup in [10] does not consider polarization reduction as an objective.
@@ 5 Conclusion
Our aim is to improve the understanding of real-world phenomena using a simplified yet insightful model. We designed a recommender system that simultaneously maximizes user engagement and mitigates polarization. Our recommender system solely relies on clicks and does not require any prior knowledge about opinion dynamics and users’ clicking behavior. We provided theoretical optimality and closed-loop guarantees for the resulting recommender-social network interconnection. Finally, our simulations demonstrated that our recommender performs favorably against other approaches that do not leverage information about users’ interconnections. We provide evidence that polarization risk should be considered at the recommendation level.
Future research directions include relaxing the smoothness assumptions on the clicking behaviour and incorporate other interest attractors towards recommendation other than confirmation bias, e.g. repulsion. Finally, we aim for our network-centered perspective to enhance the existing literature on opinion polarization caused by algorithmic systems and inspire effective mitigation strategies.
@@ 5.2 Proof of Lemma 2
$|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|\leq$ $3\underset{y_{i}\in\mathcal{X}_{y,i},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y,p)-\hat{\beta}_{i}(y,p)|+$
@@ 5.4 Hyper-parameter Tuning
Proof 5.14.
The proof is given in [43, Lemma 2].

### v2

@@ 4.1 Opinion Dynamics and Clicking Behaviour
As a proof of concept, we consider a platform with $n=15$ users. We generate the adjacency matrix $A$ with some random entries zero and other entries sampled from $\mathcal{U}[0,1]$ [^2] This ensures that the non-zero weights in the graph are not insignificant. such that $A\mathbf{1}_{n}\leq\mathbf{1}_{n}$. Thus, the resulting graph need not be strongly connected. We generate the initial opinions as $x^{0}\sim\mathcal{U}[-1,1]^{n}$. Moreover, we set $\Gamma_{d}=0.5\mathrm{diag}[|x^{0}|]$. This choice of $\Gamma_{d}$ indicates that the users who hold initial extreme opinions tend to be more stubborn. The external influence $d$ is modelled as a prejudice term defined as the initial opinion $x^{0}$. Further, $\Gamma_{p}$ is drawn from $\mathcal{D}_{[10^{-2},0.5]}^{n}$ so that all users are affected by the recommendations. To incorporate diversity in clicking behaviours, we randomly assign $8$ users to follow clicking behaviour $C_{A}$ and the remaining ones follow $C_{B}$.
@@ 4.2 Experimental Setting and Hyperparameter Selection
In the online phase, we choose $\sigma_{\text{pe}}=0.07$ in (16) for the first $N/5$ time steps to facilitate sensitivity estimation [^3] Sensitivity estimation can be halted after convergence on its entries. and network learning, and $\sigma_{\text{pe}}=0$ afterwards to favor accurate recommendations. In the Kalman filter (3.2.1), we initialize ^$\hat{\ell}^{0}=0.5{\rm vec}(I_{n})$ and $\Sigma^{0}=I_{n^{2}}$. For gradient estimation in (14)–(15), we set the smoothing parameter $\mu=0.1$. While tuning the process noise standard deviation, we set the tuning parameter $M$ in (21) as $M=10$. Further, we start with neutral recommendations, i.e., $p^{0}=0_{n}$, and consider a periodic trigger with period $T=60$. For the polarization cost $\varphi^{\rm pol}$ in (4), we set $\epsilon_{1}=-0.5$ and $\epsilon_{2}=0.5$. All the simulations are conducted for $N=2.5\cdot 10^{4}$ time instances and over $50$ Monte-Carlo trials. Since clicks are stochastic, the sequence of clicks varies across each trial. Moreover, the dither signal $w_{\rm pe}$ sequence varies across each trial, thus leading to different sensitivity estimates.
@@ 4.3 Architecture Comparisons
In Table 2, we summarize all the methods used for comparison and their attributes. In particular, for the oracle (method $M_{1}$), we employ the standard projected-gradient controller (6), wherein opinions can be directly measured, and the sensitivity and users’ clicking behavior are both known.
@@
Figure 3 shows that convergence is empirically established in all the methods through the fixed-point residual mapping norm $\lVert\mathcal{G}(p^{k})\rVert^{2}$. In particular, for methods that involve sensitivity estimation (i.e., $M_{2}-M_{4}$), we observe a higher fluctuation for the first $N/5$ time instances due to the effect of the dither signal $\sigma_{\text{pe}}$. Our proposed method offers slightly inferior performances compared to the others, which is unsurprising as it is completely data-driven.
@@
Figure 4: Evolution of the sensitivity estimation error in methods $M_{2}-M_{4}$. The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
@@
4.4 Impact of Network-aware Recommendations
@@
In this case study, we showcase the impact of providing network-aware recommendations compared to individual, decoupled recommendations. To achieve this, we compare our approach with two network-agnostic methods: (i) the recommender system described in [9], and (ii) a naive version of our OFO-based recommender design.
@@
The design in [9] provides recommendations with exclusively extreme positions (i.e., $p_{i}=\pm 1$) to each user independently. The position is calculated based on the user’s individual clicking history. Specifically, if a user more frequently clicks on $+1$ (or $-1$) positions, the system will provide $+1$ (or $-1$) positions with high probability. Conversely, if the user shows an equal likelihood of clicking on both $+1$ and $-1$ positions, the recommendations will be random. It’s important to note that the comparison with the method in [9] is not entirely equitable in terms of methodology: their approach relies solely on extreme positions and does not target polarization reduction as an objective. However, we include this comparison because, to the best of our knowledge, it is the only model-free, opinion agnostic, systems-theory-based recommender design available.
@@
The second method considered for comparison is a network-agnostic version of our recommender system design summarized in Algorithm 1. We will refer to this method as naive OFO. Specifically, for each user $i\in[n]$, we use a local ANN for opinion estimation that is trained using local positions and acceptance ratio, i.e., $\hat{x}_{i}=\hat{\beta}_{i}(y_{i},p_{i})$ rather than $\hat{\beta}_{i}(y_{i},p)$. Further, we do not carry out sensitivity estimation and instead use a constant diagonal sensitivity $\hat{H}\in\mathcal{D}_{(0,0.5]}^{n}$ in (17). The rationale behind this design choice is that if the network did not contribute to engagement maximization and/or polarization mitigation, a diagonal $\hat{H}$ would yield the same performance as our network-aware approach.
@@
We compare the performance of our recommender system with the other decoupled methods in terms of final mean CTR $J:=\sum_{k=N-T+1}^{N}\lVert c^{k}\rVert_{1}/(nT)$ and final polarization cost $\varphi^{\rm pol}(x^{N})$ in Figure 7 and the fixed point residual mapping in Figure 8. In Figure 7, we observe that our proposed algorithm achieves a slightly higher ($\sim 1.3\%$) user engagement compared to its naive implementation and results in a lower polarization cost ($\sim 9.3\%$). Thus, network-aware recommendations effectively reduce polarization while maintaining high user engagement. This performance improvement is further illustrated in Figure 8, which depicts the temporal evolution of the fixed-point residuals. This result is unsurprising since polarization is a global network metric, whereas engagement maximization can be addressed individually by tailoring recommended content for each user.
@@
Note that the recommender design in [9] employs only extreme recommendations, naturally leading to higher polarization costs, as shown in Figure 7(b). Furthermore, it results in lower user engagement compared to both our algorithm and the naive method. This is because some users in the network are assigned clicking behavior $C_{B}$, for which extreme recommendations do not necessarily increase engagement.
@@ 5 Conclusion
We introduced a novel recommender system algorithm based on online feedback optimization that balances the trade-off between maximizing user engagement (CTR) and minimizing polarization. By integrating auxiliary estimation blocks, our approach overcomes the limitations of existing OFO designs, which require unavailable real-time data, and instead relies solely on click data. We established theoretical guarantees for closed-loop stability and validated our design through numerical simulations using an extended FJ model. Our results demonstrated the impact of network-awareness in reducing polarization while maintaining high engagement.
Future compelling research directions include relaxing the smoothness assumptions for clicking behavior and exploring alternative interest drivers, such as repulsion, beyond confirmation bias. Parameter estimation for clicking behavior could also be refined by using a grey-box model instead of a purely black-box approach. Additionally, investigating the impact of different graph topologies on polarization and developing distributed algorithms for more efficient sensitivity estimation in large-scale networks would be valuable. Lastly, we hope our network-centered approach contributes to the literature on mitigating opinion polarization through algorithmic systems.
@@ 5.2 Proof of Lemma 2
$|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|\leq$ $3\underset{y_{i}\in\mathcal{X}_{y,i},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|+$
@@ 5.4 Hyper-parameter Tuning
Proof 5.15.
The proof is given in [49, Lemma 2].